# Memory Self-Regeneration: Uncovering Hidden Knowledge in Unlearned Models

### v1

@@ 5 Conclusion
Limitations While MemoRa demonstrates strong ability to rapidly recover knowledge in the case of short-term forgetting, its effectiveness is notably reduced for long-term forgetting. This highlights that our strategy excels at shallow memory restoration but struggles when the erased concept has been more deeply replaced within the model.
@@
Appendix
In the supplementary materials, we provide additional experiments. Section A details the additional implementation, training, and evaluation configurations. Appendix B presents further qualitative results for nudity, objects, explicit content, and artistic style. For all experiments, we used the NVIDIA RTX 4090 GPU. The source code will be available on GitHub after review.
@@ Appendix A Training and Experimental Setup
Ewaluation Setup To investigate the phenomenon of Memory-Self Regeneration of memory, modern methods of unlearning in diffusion models were used for testing. Each unlearning technique has a different way of forgetting a specific concept. We investigate whether each of them only superficially unlearned Stable Diffusion. Additionally, we ask whether this unlearning has already become permanently embedded in long-term memory.
@@ B.1 Nudity
Table 7: Prompts from the I2P dataset to generate nudity-style images. Images generated using these prompts are shown in Fig. 12, 13 and 14
@@
It is worth mentioning in this case that the effectiveness of the attack measure is not as reliable. Here, an attack is deemed successful if even a single feature associated with illegal content is detected at a low threshold. Consequently, this measure only indicates the recovery of individual features and it does not adequately capture the overall concept of nudity. Fig. 15 presents the distribution of nudity classes, in which the situation described earlier occurs, where MemoRa generates more features related to the nudity concept, where UnlearnDiffAtk looks for one feature to consider the attack successful.
@@
Figure 17 presents cosine similarity histograms comparing the unlearned model (SalUn) and SalUn+MemoRa against SD v1.4 for the same prompts. For I2P, the unlearned model shows low similarity, indicating effective forgetting. After applying the MemoRa Memory Self-Regeneration strategy, the histogram shifts noticeably toward higher values, suggesting that the model’s outputs are closer to those of the base SD v1.4. Importantly, a similar trend is observed for safe prompts from the MS-COCO dataset, demonstrating that the recall strategy increases similarity without compromising general generation quality.
@@ B.2 Objects
This section provides supplementary visual and quantitative comparisons related to the unlearning and MSR of objects, including a church, parachute, and garbage truck, see Fig. 18. We show that with the help of our MemoRa strategy, the restoration of the objects was possible, and visually they are similar to the original SD.
Restoring objects presents an additional challenge, as objects can often be confused with similar ones. UnlearnDiffAtk illustrates this situation well in Fig. 19. Often, to accurately recover a church, it’s not enough to generate the building itself. The same situation applies to restoration, for example, of a garbage truck, see Fig. 23. A model that has successfully reinstated knowledge about a garbage truck must accurately draw this object so that the classifier does not confuse it with a truck or a car during evaluation.
Tab. 8 shows the numerical results for MemoRa on the memory recall task for the concept church. Significant increase in recall is observed for the SalUn, ESD, SPM, ED techniques. Fig. 19 shows a visual comparison of our strategy with the UnlearnDiffAtk method. The first part of the table (top) presents the results for the ESD method, whose unlearning was only shallow. In the second part (bottom), the SalUn method unlearned the object more strongly, but the MemoRa strategy was also able to quickly restore the church.
@@
Fig. 21 presents a comparison of the UnlearnDiffAtk method with our MemoRa strategy, but this time for the parachute. Additional images are shown in Fig. 22.
Fig. 23 illustrates the retrieval of the garbage truck concept by MemoRa compared to prompt attacks. This object proved to be one of the most difficult concepts to restore (see Tab 9). Despite this, MemoRa restored knowledge for each method (except SH). Techniques that consolidated forgotten concepts into long-term memory include SPM, AdvUnlearn, ED, and SH. SalUn, as before, superficially unlearned the garbage truck (2% success rate), as our strategy significantly restored its knowledge (52%). For the SH method, neither our strategy nor the prompt attack method restored any knowledge, indicating deeply ingrained forgetting. The only advantage is the lower FID obtained with MemoRa (a drop from 71.17% to 55.59%), which emphasizes a slight recovery of the model to its initial state.
UnlearnDiffAtk also encountered problems in restoring this object. A potential method to obtain better results is to combine UnlearnDiffAtk + MemoRa. For this configuration, for almost all techniques, the numerical results fluctuate around 100%. Additional images are shown in Fig. 24.
@@ B.3 Styles
The ViT-base model, fine-tuned on the WikiArt dataset, was used as a classifier to evaluate the attacks. The classifier exhibits high uncertainty in classification, so similarly to Zhang et al. (2024b), we consider the top-3 score metric in Table 10.

### v2

@@ 6 Conclusion
Limitations: While MemoRa demonstrates strong ability to rapidly recover knowledge in the case of short-term forgetting, its effectiveness is notably reduced for long-term forgetting. This highlights that our strategy excels at shallow memory restoration but struggles when the erased concept has been more deeply replaced within the model.
Appendix
In the supplementary materials, we provide additional experiments. Section A details the additional implementation, training, and evaluation configurations. Appendix B presents further qualitative results for nudity, objects, explicit content, and artistic style. For all experiments, we used the NVIDIA RTX 4090 GPU (SD 1.4) and DGX H100 GPU (FLUX). The source code will be available on GitHub after review.
@@ Appendix A Training and Experimental Setup
Evaluation Setup To investigate the phenomenon of Memory-Self Regeneration of memory, modern methods of unlearning in diffusion models were used for testing. Each unlearning technique has a different way of forgetting a specific concept. We investigate whether each of them only superficially unlearned Stable Diffusion. Additionally, we ask whether this unlearning has already become permanently embedded in long-term memory.
@@
Trade-off Between Re-Learning and Utility During inference, the LoRA adapter for all tasks causes the FID metric to increase compared to the unlearned model, while the CLIP metric remains high, as shown in Tab. 1 and Tab. 12. To address this issue, we employ the Autoguidance technique (Kasymov et al., 2024; Karras et al., 2024), which involves using a combination (interpolation) of responses from both models during generation, rather than relying on just one model.
@@
Often, a model that has undergone unlearning may forget multiple concepts. We demonstrate the use of multi-MemoRa on the task of relearning well-known celebrities. Combining adapters and recalling multiple items for our strategy is not difficult. The images in Fig. 13 demonstrate that the MACE model only shallowly unlearned the selected celebrities, as Multi-MemoRa effectively restored knowledge about famous people.
@@ B.1 Nudity
Table 11: Prompts from the I2P dataset to generate nudity-style images. Images generated using these prompts are shown in Fig. 14, 15 and 16
@@
It is worth mentioning in this case that the effectiveness of the attack measure is not as reliable. Here, an attack is deemed successful if even a single feature associated with illegal content is detected at a low threshold. Consequently, this measure only indicates the recovery of individual features and it does not adequately capture the overall concept of nudity. Fig. 11 presents a comparison between concept recovery for UnlearnDiffAtk and MemoRa. Both strategies add the forbidden concept, but the images for MemoRa contain more nudity features in individual photos. In particular, the Belly, Female Genitals, and Male Breast. UnlearnDiffatk aims to generate at least one forbidden feature, while LoRA applies the nudity concept overall, creating photos more similar to the original ones. Fig. 17 presents the distribution of nudity classes, in which the situation described earlier occurs, where MemoRa generates more features related to the nudity concept, where UnlearnDiffAtk looks for one feature to consider the attack successful.
@@
Fig.19 presents cosine similarity histograms comparing the unlearned model (SalUn) and SalUn+MemoRa against SD v1.4 for the same prompts. For I2P, the unlearned model shows low similarity, indicating effective forgetting. After applying the MemoRa Memory Self-Regeneration strategy, the histogram shifts noticeably toward higher values, suggesting that the model’s outputs are closer to those of the base SD v1.4. Importantly, a similar trend is observed for safe prompts from the MS-COCO dataset, demonstrating that the recall strategy increases similarity without compromising general generation quality.
@@ B.2 Objects
Dataset Metrics SD v1.4 FMN SPM ESD
base unlearn MemoRa unlearn MemoRa unlearn MemoRa
CLIP $(\uparrow)$ 31.08 30.69 31.21 31.07 31.34 29.95 30.84
Dataset Metrics AdvUnlearn SalUn ED SH
unlearn MemoRa unlearn MemoRa unlearn MemoRa unlearn MemoRa
CLIP $(\uparrow)$ 30.56 30.44 31.13 31.16 30.88 31.00 27.72 27.68
Metric FLUX. 1 [dev] ESD UCE
base unlearn MemoRa unlearn MemoRa
No Attack $(\uparrow)$ 100% 0% 61% 35% 59%
Table 13: Evaluation of Parachute Concept Memory Recovery for FLUX.1 [dev] model. The values in per cents indicate the amount of detected parachutes in images generated using the same prompts and seeds for different models.
This section provides supplementary visual and quantitative comparisons related to the unlearning and MSR of objects (Fig. 20), including a church (Fig. 21), parachute (Fig. 22, 23), and garbage truck (Fig. 24). We show that with the help of our MemoRa strategy, the restoration of the objects is possible.
Tab. 12 referring to unlearning-relearning concept parachute. We can observe the weakest unlearning for FMN, SPM. Applying MemoRa sheds light on the actual forgetting, revealing that some methods achieve high recall rates in an instant, these are: FMN, SPM, ESD, SalUn, and ED. Forgotten knowledge about concept can be recovered to approximately 85% from just a few percent and even 100% after using additional attacks. In summary, more effective unlearning does not always translate into greater resistance to knowledge retrieval. Fig. 25 presents a comparison of the UnlearnDiffAtk method with our MemoRa strategy for the parachute.
In Fig. 22 the results of using MemoRa to restore the church are presented, where the effectiveness of the strategy is noticeable. MemoRa can recreate the original photo even for methods that significantly changed the trajectory for the unlearned concept. Restoring objects presents an additional challenge, as objects can often be confused with similar ones. UnlearnDiffAtk illustrates this situation well in Fig. 26. The first part of the table (top) presents the results for the ESD method, whose unlearning was only shallow. In the second part (bottom), the SalUn method unlearned the object more strongly, but the MemoRa strategy was also able to quickly restore the church. Tab. 14 shows the numerical results for MemoRa on the memory recall task for the concept church. Significant increase in recall is observed for the SalUn, ESD, SPM, ED techniques.
Similar observation can be observed in task of restoration of a garbage truck concept, see Fig. 27. A model that has successfully reinstated knowledge about a garbage truck must accurately draw this object so that the classifier does not confuse it with a truck or a car during evaluation.
Tab. 13 illustrates the effect of recalling knowledge about parachutes. Similar to SD v1.4, ESD demonstrates a shallow level of unlearning. Fig. 4 displays the visual results of the MemoRa effect. Interestingly, the ESD technique significantly transformed the photographs, creating different scenes. In contrast, UCE subtly removed only the parachute object. For the nudity concept, the numerical results are presented in Tab. 6, where the unlearned model ultimately returned very close to the baseline state. More visual results are shown in Fig. 2.
@@
Fig. 27 illustrates the retrieval of the garbage truck concept by MemoRa compared to prompt attacks. This object proved to be one of the most difficult concepts to restore (see Tab 15). Despite this, MemoRa restored knowledge for each method (except SH). Techniques include SPM, AdvUnlearn, ED, and SH. SalUn, as before, superficially unlearned the garbage truck (2% success rate), as our strategy significantly restored its knowledge (52%). For the SH method, neither our strategy nor the prompt attack method restored any knowledge, indicating deeply ingrained forgetting. The only advantage is the lower FID obtained with MemoRa (a drop from 71.17% to 55.59%), which emphasizes a slight recovery of the model to its initial state.
UnlearnDiffAtk also encountered problems in restoring this object. A potential method to obtain better results is to combine UnlearnDiffAtk + MemoRa. For this configuration, for almost all techniques, the numerical results fluctuate around 100%.
@@ B.3 Styles
Style Relearning
For style evaluation, prompts were used according to the setup described in (Gandikota et al., 2023).
The ViT-base model, fine-tuned on the WikiArt dataset, was used as a classifier to evaluate the attacks. The classifier exhibits high uncertainty in classification, so similarly to Zhang et al. (2024b), we consider the top-3 score metric in Table 16. However, the classifier may struggle to correctly classify such paintings, as the paintings are very intense and slightly deviate from the original.
Fig. 28 presents a comparison between images before and after the MemoRa strategy. Images generated after applying the reminder strategy demonstrate the distinctive style of Vincent Van Gogh. MemoRa accurately reproduced the Starry Night scene, incorporating strong brushstrokes, swirls in the sky, and a contrast of blues and yellows. Therefore, it seems that the models do not long-term forget about this distinctive style.